# One-Prompt to Segment All Medical Images

### v3

@@ Abstract
Large foundation models, known for their strong zero-shot generalization, have excelled in visual and language applications. However, applying them to medical image segmentation, a domain with diverse imaging types and target labels, remains an open challenge. Current approaches, such as adapting interactive segmentation models like Segment Anything Model (SAM), require user prompts for each sample during inference. Alternatively, transfer learning methods like few/one-shot models demand labeled samples, leading to high costs. This paper introduces a new paradigm toward the universal medical image segmentation, termed ’One-Prompt Segmentation.’ One-Prompt Segmentation combines the strengths of one-shot and interactive methods. In the inference stage, with just one prompted sample, it can adeptly handle the unseen task in a single forward pass. We train One-Prompt Model on 64 open-source medical datasets, accompanied by the collection of over 3,000 clinician-labeled prompts. Tested on 14 previously unseen tasks, the One-Prompt Model showcases superior zero-shot segmentation capabilities, outperforming a wide range of related methods. The code and annotated data will be publicly released.
@@ 1 Introduction
Large foundation models pre-trained on large-scale datasets, are transforming the landscape with powerful zero-shot capabilities [54, 27]. These foundational models showcase an impressive ability in adapting to the tasks not seen during training. A standout example is the Segment Anything Model (SAM) [27], which has gained great success for the zero-shot image segmentation. The strength of SAM lies in its interactive segmentation paradigm: the model segments the target following the user-given prompts, such as a point, a bounding box (BBox), or free text-like descriptions.
@@
In this paper, we introduce a new paradigm for the universal medical image segmentation, called One-Prompt Medical Image Segmentation. This method combines the strengths of both one-shot and interactive models to meet the real clinical requirements. Specifically, given an unseen task, the user only needs to provide one prompted sample to the trained model, then it can perform well at this new task without any retraining or fine-tuning, even for tasks significantly different from those encountered during training. An illustration is shown in Fig. 1. In essence, the one-prompt model can adapt to any unseen medical image segmentation task by seeing just one prompted sample through one forward pass.
The success of the One-Prompt Model is driven by three key factors. First, we propose novel One-Prompt Former modules as the model decoder. Such design helps to efficiently integrate the prompted template feature into the process of query image segmentation. Secondly, we gather a large-scale data collection comprising 78 open-source datasets covering diverse biomedical domains. Our model is trained on 64 datasets, with clinicians prompting 10% of the data. Moreover, to further enhance the clinical utility, we offer four different prompt types, which are Click, BBox, Doodle, and SegLab. The Click and BBox work the same as those in SAM [27]. Doodle allows the users to freely draw on the image, helpful for prompting irregular organs like the pancreas or lymph glands. SegLab allows the user to provide the segmentation mask as the prompt for representing more detailed tissues like vessels.
@@
  • We gather a large-scale collection comprising 78 open-source datasets to train and test the model. We also annotated 10% of the samples with clinician-given prompts, which will be released for the further research purposes.
@@ 2.2 Model
The One-Prompt Model comprises a standard CNN-based image encoder and a sequence of One-Prompt Former as the decoder, illustrated in Fig. 2 (a). The model takes three inputs: the query image $x_{q}$, the template image $x_{t}$, and the prompt of the template image $p_{t}$, and subsequently predicts the segmentation of the query denoted as $y_{q}$. The multi-scale features of the encoder and decoder are skip-connected.
Encoder The image encoder follows the standard implementation of a generic UNet [24] encoder that sequentially abstracts the feature to 1/4, 1/8, 1/16, and 1/32 downsampled ones. The query sample $x_{q}$ and the template sample $x_{t}$ will both go through the same encoder to get the feature $f_{q}$ and $f_{t}$.
@@ 2.3 Training and Loss
We divided our One-Prompt dataset into 64 datasets for model training and 14 datasets for testing. Each training dataset is further split into three parts: a prompted template split (10% of samples), a training split (70% of samples), and a validation split (20% of samples). Similarly, each test dataset has a test split (90% of samples) and a prompted template split (10% of samples). Human users prompt each sample in the template split, for both training and testing. Model training is performed on the template and training splits across all datasets. In each iteration, we randomly pick one prompted template from the template set of the same dataset with the query image. Training is collectively conducted across the 64 datasets. Our final loss is a simple sum of cross-entropy loss and dice loss.
@@ 3.1 Data Source
In order to construct a foundation model with high generalization on the unseen tasks, we train our model on large-scale and diverse medical images consisted by online open-access datasets. Our data source is constructed from 78 datasets encompassing diverse medical domains and imaging modalities. The dataset covers a wide array of organs, such as lungs [48, 49, 50], eyes [16, 23, 39, 40], brain [5, 17, 22, 28, 29], and abdominal [25, 6, 20, 26, 30, 31, 32, 33, 35, 38, 44, 50]. Supplemental Section A provides a detailed list of One-Prompt datasets.
@@ 3.2 Prompt Annotation
A team of clinicians prompt about 10% of our samples on each training dataset. These samples are meticulously selected by experienced annotators to ensure diversity and comprehensiveness. The clinicians involved in prompting come from diverse backgrounds, including cardiologists, dermatologists, gastroenterologists, neurologists, oncologists, pulmonologists, rheumatologists, endocrinologists, and ophthalmologists. They are encouraged to choose datasets aligned with their expertise during the prompting process.
In this process, all four prompt types are available for each sample, and clinicians are encouraged to use the most convenient prompt tool for the given targets. The clinicians employ a browser-based interactive segmentation tool to prompt images. Upon prompting, ground-truth masks immediately appear on the images based on the given prompts. Clinicians have the flexibility to refine their prompts, but adjustments are suggested only if they feel their initial prompt was incorrect. Our prompt-based segmentation operates in real-time directly within a browser. Notably, we do not set strict constraints on prompt quality Clinicians are encouraged to prompt images in their most natural way, with a suggested time limit of no more than 5 seconds for each image. The prompting details on the test set can be found in Section 4.1.
@@ 4 Experiments
In this section, we present task generalization experiments with One-Prompt Model. We divide out 14 tasks in our available datasets, which differ significantly in term of image modalities and target structures, as our held-out test set. This set comprises 8 MICCAI2023 Challenge tasks, encompassing various anatomies including kidney tumor [21], liver tumor [43], breast cancer [1], nasopharynx cancer [4], vestibular schwannoma [2], mediastinal lymph node [45], cerebral artery [11], and inferior alveolar nerve [7]. The other 6 tasks including the segmentation of white blood cell [58], optic cup [16], mandible [3], coronary artery [52], pancreas [46], and retinal blood vessel [23]. Among them, three datasets [46, 23, 1] contain anatomies represented in the training set, while others introduce anatomies not covered in training.
We assess the model performance on each test dataset using a specific prompt type, informed by the observation that users tend to favor specific prompts for particular tasks. For instance, in pancreas segmentation, 93% of samples are prompted by Doodle, while in nasopharynx cancer gross tumor segmentation, 87% of samples are prompted by BBox. Such a comparison setting allows us to capture the dominant use case for each dataset. Because of the space constraint, we provide our implementation and data processing details in the appendix.
@@ 4.1 Human-User Prompted Evaluation
For the evaluation, we involved human users to simulate real-world interactions for prompt-based segmentation. We assigned 15 users to prompt a total of 10% of the test images. The users comprised 5 regular individuals with a clear understanding of the task but no clinical background, 7 junior clinicians, and 3 senior clinicians. This aims to simulate real-world prompting scenarios such as clinical education or semi-automatic annotation.
@@ 4.3 Interactive Segmentation Capability
Task Interactive segmentation models achieve zero-shot generalization by prompting each of the test sample. When we offer the One-Prompt Model with the same query image and prompted template image, the model degrades to a standard interactive segmentation model. We compare this setting with other interactive segmentation models, including vanilla SAM [27], SAM-U [14], VMN [59], iSegFormer[34], MedSAM [37], MSA [56], and SAM-Med2D [13]. Except vanilla SAM, all models are trained on the same dataset as ours. Since most of these models only accept Click and BBox prompts, Doodle and SegLab prompt settings are not included in this comparison. Since all these models need the prompt on each input image, we simulate the oracle prompts (details in Section 4.5: Effect of prompt quality & types in the inference) over the images if needed. It is worth noting that we did not re-train One-Prompt Model on the simulated prompts: we use the same trained One-Prompt Model as that in the last section, but only offered the simulated prompts in testing for the possible of comparison.
@@
Results We present a quantitative comparison with the interactive segmentation methods in Fig. 4. We can see in the figure that One-Prompt Model outperforms all other interactive competitors by a significant margin. These results demonstrate that One-Prompt Model could perform as well when the query image itself is prompted, despite not being intentionally trained for this specific setting. By training under our more challenging setting (predicting 70% of samples with only 10% prompts), the One-Prompt Model demonstrates superior capability compared to interactive models.
@@
4.4 Zero-shot Capability
Data We use 11 unseen tasks in the held-out test set to verify the zero-shot transfer ability of the models. The target anatomies of these tasks are not appeared in the training. Detailed datasets are shown in Table. 1

### v5

@@ Abstract
Large foundation models, known for their strong zero-shot generalization, have excelled in visual and language applications. However, applying them to medical image segmentation, a domain with diverse imaging types and target labels, remains an open challenge. Current approaches, such as adapting interactive segmentation models like Segment Anything Model (SAM), require user prompts for each sample during inference. Alternatively, transfer learning methods like few/one-shot models demand labeled samples, leading to high costs. This paper introduces a new paradigm toward the universal medical image segmentation, termed ’One-Prompt Segmentation.’ One-Prompt Segmentation combines the strengths of one-shot and interactive methods. In the inference stage, with just one prompted sample, it can adeptly handle the unseen task in a single forward pass. We train One-Prompt Model on 64 open-source medical datasets, accompanied by the collection of over 3,000 clinician-labeled prompts. Tested on 14 previously unseen datasets, the One-Prompt Model showcases superior zero-shot segmentation capabilities, outperforming a wide range of related methods. The code and data is released as https://github.com/KidsWithTokens/one-prompt.
Apology: We sincerely apologize to Jiayuan Zhu and Yueming Jin for their absence from the authorship list, despite their significant contributions to this project. Unfortunately, due to missing the CVPR registration deadline, we were unable to include them. We deeply appreciate their dedication, expertise, and insights, which have been instrumental in bringing this work to fruition.
@@ 1 Introduction
Large foundation models pre-trained on large-scale datasets, are transforming the landscape with powerful zero-shot capabilities [56, 27, 55, 54]. These foundational models showcase an impressive ability in adapting to the tasks not seen during training. A standout example is the Segment Anything Model (SAM) [27], which has gained great success for the zero-shot image segmentation. The strength of SAM lies in its interactive segmentation paradigm: the model segments the target following the user-given prompts, such as a point, a bounding box (BBox), or free text-like descriptions.
@@
In this paper, we introduce a new paradigm for the universal medical image segmentation, called One-Prompt Medical Image Segmentation. This method combines the strengths of both one-shot and interactive models to meet the real clinical requirements. Specifically, given an unseen task, the user only needs to provide one prompted sample to the trained model, then it can perform well at this new task without any retraining or fine-tuning, even for tasks significantly different from those encountered during training. An illustration is shown in Fig. 1.
The success of the One-Prompt Model is driven by three key factors. First, we propose novel One-Prompt Former modules as the model decoder. Such design helps to efficiently integrate the prompted template feature into the process of query image segmentation. Secondly, we gather a large-scale data collection comprising 78 open-source datasets covering diverse biomedical domains. Our model is trained on 64 datasets, with clinicians prompting a part of the data. Moreover, to further enhance the clinical utility, we offer four different prompt types, which are Click, BBox, Doodle, and SegLab. The Click and BBox work the same as those in SAM [27]. Doodle allows the users to freely draw on the image, helpful for prompting irregular organs like the pancreas or lymph glands. SegLab allows the user to provide the segmentation mask as the prompt for representing more detailed tissues like vessels.
@@
  • We gather a large-scale collection comprising 78 open-source datasets to train and test the model, and also annotated over 3000 of the samples with clinician-given prompts.
@@ 2.2 Model
The One-Prompt Model comprises an image encoder and a sequence of One-Prompt Former as the decoder, illustrated in Fig. 2 (a). The model takes three inputs: the query image $x_{q}$, the template image $x_{t}$, and the prompt of the template image $p_{t}$, and subsequently predicts the segmentation of the query denoted as $y_{q}$. The multi-scale features of the encoder and decoder are skip-connected.
Encoder The image encoder can be CNN based [24] or ViT based [12]. We show a CNN based encoder in the figure for the simple illustration. The query sample $x_{q}$ and the template sample $x_{t}$ will both go through the same encoder to get the feature $f_{q}$ and $f_{t}$.
@@ 2.3 Training and Loss
We divided our One-Prompt dataset into 64 datasets for model training and 14 datasets for testing. Each training dataset is further split into three parts: a prompted template split, a training split, and a validation split. Similarly, each test dataset has a test split and a prompted template split. Human users prompt each sample in the template split, for both training and testing. Model training is performed on the template and training splits across all datasets. In each iteration, we randomly pick one prompted template from the template set of the same dataset with the query image. Training is collectively conducted across the 64 datasets. Our final loss is a simple sum of cross-entropy loss and dice loss.
@@ 3.1 Data Source
In order to construct a foundation model with high generalization on the unseen tasks, we train our model on large-scale and diverse medical images consisted by online open-access datasets. Our data source is constructed from 78 datasets encompassing diverse medical domains and imaging modalities. The dataset covers a wide array of organs, such as lungs [48, 49, 50], eyes [16, 23, 39, 40], brain [5, 17, 22, 28, 29], and abdominal [25, 6, 20, 26, 30, 31, 32, 33, 35, 38, 44, 50]. A detailed list of One-Prompt datasets is released with our code.
@@ 3.2 Prompt Annotation
A team of clinicians prompt over 3000 samples across over all collected dataset. These samples are meticulously selected by experienced annotators to ensure diversity and comprehensiveness. The clinicians involved in prompting come from diverse backgrounds, including cardiologists, dermatologists, gastroenterologists, neurologists, oncologists, pulmonologists, rheumatologists, endocrinologists, and ophthalmologists. They are encouraged to choose datasets aligned with their expertise during the prompting process.
In this process, all four prompt types are available for each sample, and clinicians are encouraged to use the most convenient prompt tool for the given targets. The clinicians employ a browser-based interactive segmentation tool to prompt images. Upon prompting, ground-truth masks immediately appear on the images based on the given prompts. Clinicians have the flexibility to refine their prompts, but adjustments are suggested only if they feel their initial prompt was incorrect. Our prompt-based segmentation operates in real-time directly within a browser. Notably, we do not set strict constraints on prompt quality. Clinicians are encouraged to prompt images in their most natural way, with a suggested time limit of no more than 5 seconds for each image. The prompting details on the test set can be found in Section 4.1.
@@ 4 Experiments
In this section, we present task generalization experiments with One-Prompt Model. We divide out 14 tasks in our available datasets, which differ significantly in term of image modalities and target structures, as our held-out test set. This set comprises 8 MICCAI2023 Challenge tasks, encompassing various anatomies including kidney tumor [21], liver tumor [43], breast cancer [1], nasopharynx cancer [4], vestibular schwannoma [2], mediastinal lymph node [45], cerebral artery [11], and inferior alveolar nerve [7]. The other 6 tasks including the segmentation of white blood cell [60], optic cup [16], mandible [3], coronary artery [52], pancreas [46], and retinal blood vessel [23]. We assess the model performance on each test dataset using a specific prompt type, informed by the observation that users tend to favor specific prompts for particular tasks. We provide our implementation and data processing details in the appendix.
@@ 4.1 Human-User Prompted Evaluation
For the evaluation, we involved human users to simulate real-world interactions for prompt-based segmentation. We assigned 15 users to prompt about 10% of the test images. The users comprised 5 regular individuals with a clear understanding of the task but no clinical background, 7 junior clinicians, and 3 senior clinicians. This aims to simulate real-world prompting scenarios such as clinical education or semi-automatic annotation.
@@ 4.3 Interactive Segmentation Capability
Task
Interactive segmentation models achieve zero-shot generalization by prompting each of the test sample. When we offer the One-Prompt Model with the same query image and prompted template image, the model degrades to a standard interactive segmentation model. We compare this setting with other interactive segmentation models, including vanilla SAM [27], SAM-U [14], VMN [61], iSegFormer[34], MedSAM [37], MSA [58], and SAM-Med2D [13]. Except vanilla SAM, all models are trained on the same dataset as ours. Since most of these models only accept Click and BBox prompts, Doodle and SegLab prompt settings are not included in this comparison. Since all these models need the prompt on each input image, we simulate the oracle prompts (details in Section 4.5: Effect of prompt quality & types in the inference) over the images if needed. It is worth noting that we did not re-train One-Prompt Model on the simulated prompts: we use the same trained One-Prompt Model as that in the last section, but only offered the simulated prompts in testing for the possible of comparison.
@@
Results. We present a quantitative comparison with the interactive segmentation methods in Fig. 4. We can see in the figure that One-Prompt Model outperforms all other interactive competitors by a significant margin. These results demonstrate that One-Prompt Model could perform as well when the query image itself is prompted, despite not being intentionally trained for this specific setting. By training under our more challenging setting, the One-Prompt Model demonstrates superior capability compared to interactive models.
@@
4.4 Zero-shot Capability
Data We use 11 unseen datasets in the held-out test set to verify the zero-shot transfer ability of the models. Detailed datasets are shown in Table. 1